OpenStudy (anonymous):

If t is a prime number, and x is a factor of 12, then t/x could be all of the following EXCEPT A) 1/12 B) 1/4 C) 1/2 D) 1 E) 2

OpenStudy (anonymous):

Factor of 12: 1,2,3,4,6,12 = X Prime numbers: 1,2,3,5,7,11,13,17,19, etc. = T A) 1/12 =YES B) 1 / 4 = YES C) 1/2 = YES D) 1/1= YES E) HOW?

OpenStudy (anonymous):

what are the factors of 12? 1,2,3,4,6,12 so \(x={1,2,3,4,6,12}\) what are some prime numbers? 1,3,5,7,9,11... so \(t={1,2,3,5,7,}\) now lets put \(t\) over \(x\):\[\color{red}{1\over1},\color{red}{2\over1},{3\over1},{5\over1},{7\over1}\]\[\color{red}{1\over2},\color{red}{2\over2},{3\over2},{5\over2},{7\over2}\]\[{1\over3},{2\over3},{3\over3},{5\over3},{7\over3}\]\[\color{red}{1\over4},{2\over4},{3\over4},{5\over4},{7\over4}\]\[{1\over6},{2\over6},\color{red}{3\over6},{5\over6},{7\over6}\]\[\color{red}{1\over12},{2\over12},\color{red}{3\over12},{5\over12},{7\over12}\] they are all possible, so I'm not sure if maybe the question has been posted wrong? @iNeedAllDaHelp